OpenStudy (goformit100):

The volume of a cube is increasing at a rate of 9 cubic centimetres per second. How fast is the surface area increasing when the length of an edge is 10 centimetres ?

OpenStudy (goformit100):

@terenzreignz @Shannon20150 @Noemi95

terenzreignz (terenzreignz):

If we let S = the surface area, then we need to find... \[\huge \frac{dS}{dt}\] given of course that \[\Large \frac{dV}{dt}=9\]

OpenStudy (anonymous):

There he gave it to u

terenzreignz (terenzreignz):

It would be nice if we could express S in terms of V

OpenStudy (anonymous):

@terenzreignz's smart self got it gofor! (:

terenzreignz (terenzreignz):

Meanwhile, @terenzreignz ' childish self is still sleeping, while his hungry self is indulging in ice cream... :) If we let x = one edge of the cube, then \[\Large S = 6x^2\\\Large V=x^3\]

terenzreignz (terenzreignz):

\[\huge \frac{dS}{dV}=\frac{\frac{dS}{dx}}{\frac{dV}{dx}}\]

terenzreignz (terenzreignz):

\[\huge \frac{dS}{dt}=\frac{dS}{dV}\cdot\frac{dV}{dt}=\frac{\frac{dS}{dx}\times \frac{dV}{dt}}{\frac{dV}{dx}}\]

OpenStudy (goformit100):

Thankyou Sir

terenzreignz (terenzreignz):

And then, plug in x = 10

OpenStudy (goformit100):

HELP http://openstudy.com/study#/updates/517df6f6e4b0be6b54ab6738